OpenStudy (anonymous):

Solve the system by using elimination or elimination with multiplication. Write the solution as an ordered pair (a, b). a - b = -5 and a + 2b = 1

OpenStudy (anonymous):

subtract the two equations... what are you left with?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

a - b = -5 and a + 2b = 1 ( subtracting the two equation you get ) => 3b = 6 => b = 6/3 = 2 and putting value of b in any of the equation you get a = -3